Monthly Report IR Demo FEL Upgrade Project
November 2000




Project Highlights:

Good progress was made this week incorporating the configuration changes to the optical cavity design that were discussed at the October 23 semi-annual review of the FEL Upgrade.

We completed a very successful Fall FEL User run on Nov. 17.  The machine ran well and the users were pleased with their results which will be discussed at the January 18-19 LPC/FEL Users Workshop.
FEL operations produced a new average power record for sub-picosecond class 1 micron light by generating over 340 watts with third harmonic lasing.  We also attempted our first experiments with doubling (to green) and quadrupling (to UV) the 1 micron light achieving ~28% and ~5% conversion efficiencies on our initial trails with non-optimized non-linear conversion optics.

On Nov. 27th we received from one of our generous industrial collaborators a significant equipment
donation that will enhance the capabilities of the FEL User Facility:  a compact, superconducting
storage ring capable of generating x-rays  with a 1nm critical wavelength.   We are currently seeking
support from the Commonwealth of Virginia to construct  an addition onto the FEL building to house
the synchrotron.

Project Cost Performance:

The project budget for the period June 1, 2000 to Sept. 30, 2001 is $9,029K. The project through the month of November has a total of $1,941K of performance scheduled (assuming the project started at the originally planned start date of April 1, 2000).  The work performed through the end of November was $2,051, which is 23% complete vs. 22% scheduled.  The actual cost accrued through November totals $2,059K.  This results in a schedule variance of +$109k and a cost variance of $9K.  The favorable cost and schedule variance continues because of the large quantity of RF work (WBS 6) that was completed ahead of schedule and under cost.  Work in the Beam transport (WBS 9) and Optics (WBS 11) will continue to show minor negative schedule variances until we fully recover from the late project start and the engineering change request for the optical cavity design approved at the Oct. 23 rd review.

Management:

We received from our ONR contract monitor a draft report from the project semi-annual review that was held on Oct. 23.  To quote from the draft report:  "... plans for the 10 kW Upgrade are solid and the one major technical risk area which the panel identified at the kick-off review (the optical cavity) has been well addressed".

We notified two companies this month that they successfully competed for task order procurements we plan to place for engineering , design and analysis work on electron beam and optical beam transport hardware.  The project team spent two extended sessions this month discussing long range plans for the FEL, paying particular attention to next steps beyond the 10 kW device.  These discussions will be summarized for further use.

We submitted two proposals in response to the Joint Technology Office call for proposals on high energy laser technology.  Our response topics were:  (1) femto second laser damage effects; and (2) high average current beam transport.

We hosted visitors this month from NRL, DARPA and Lockheed Martin who are interested in discussing possible lithography activities with the synchrotron which is being considered for addition to the FEL Facility.

WBS 3 (Beam Physics):

Analysis of injector simulations at 135 pC has led to a recommendation for the use of sector-dipole injection line geometry in the upgrade. This will be incorporated in the Revision 1.1 release of the machine design.  Other features of the Accelerator Design Revision 1.1 include: space for use of 7-cell modules everywhere, revised magnet field roll-off values, and  accommodation of  the footprint of the adopted 32 m near-concentric optical cavity resonator.  Initial studies of the injection line geometry for this design has provided an explanation for the reinjection steering observed in the IR Demo - it is due to the use of sector dipoles in the reinjection line - and suggest this will be less significant in the Upgrade, where the bend angles are smaller.

As part of the long range planning sessions mentioned above, we developed a list of beam physics issues that need to be quantified by further modeling and/or measurements on the IR Demo prior to the planned stop of machine operations approximately one year from now for upgrade installation activities.

WBS 4 (Injector):

The second nitrogen implanted electrode for the field emission test system was initially run at 24 MV/m
(5mm gap at 120 kV) and had an average field emission current of 100pA.  This initial data is better
than the first implanted electrode which showed 145 pA at 24 MV/m.  The second test run on this  nitrogen implanted electrode achieved  30 MV/m with a base line current of around 180 pA which was higher than the first test electrode (55 pA) but still well below the limits of acceptability.  The electrode was run for 7 hours at 30 MV/m showing a higher but less active baseline current.  We installed a titanium cathode into the field emission test system and set up the system for a vacuum bake prior to more data collection.

WBS 5 (SRF):

In the cavity area, we are continuing to fabricate cavity segments, weld them up and prepare them for cavity tuning.  A vendor visit was made to  the waveguide supplier for kick off of the contract.  The vendor looks capable and well suited to supply parts.  The first article helium vessels were shipped this month  from PHPK, the vessel vendor.

WBS 6 (RF):

We continued installation and check-out activities in Zones 3 and 4:
Zone 4 We completed preliminary checkout of the HPA and its interlocks.  Software has been written to test the zone.  As soon as the LCW is restored to the FEL, RF testing will start.
Zone 3 - No direct progress.  The documentation is being updated as a result of the problems found in zone 4, so zone 3 should come together quickly once parts are available.

WBS 8 (I&C):

The optical beam position monitoring system (starting with can 6) is now operational.  The control screen displays both the position and incident power.  Efforts were put into post-run cleanup of signal and cables. The new VME boards for beam sync distribution have been stuffed, tested, and installed. Another optical BPM was placed at the end-of-line dump. The next cut at the closed loop system architecture was worked out with the optics group. This will consist of five OBPM monitors on the transport controlling 3 (X&Y) mirror cans. This will be tested during the Feb. FEL run.

The beam viewer contract is complete and work is progressing (the vendor buying parts and materials).

Meetings were held to review the parts count for magnets and instrumentation for the IR upgrade and the UV add-on. The viewers and BPMs were still consistent with the initial estimates made last summer.
Regarding trim magnet control the current approach is to combine all iron core magnets into the existing trim racks (utilizing the on/off hysteresis loop capabilities) and look for a lower cost bipolar supply for the air core magnets.  This could result in cost savings of $200k (90 air core with <$1k/ch vs. trim rack cost of ~$3k/ch). Commercial vendors are being consulted.

WBS 9 (Transport):
Dipoles
We sent out preliminary dipole drawings with a request for a budgetary price to six vendors. George Biallas walked one vendor through the package during the same trip as a Berkeley Lab magnet review.  GW (optical cavity dipole) prototype drawings are continuing to be developed and detailed.  3D magnetic models of the GW were run using three different materials for the Purcell Gap and at multiple field strengths. It was concluded that the Purcell Gap offered no improvement to the good field and would not be required to first order and contrary to our experience.  We believe the need for the Purcell Gap comes from residual field from hysteresis, changes in the field that are not modeled by existing software.  We are making the GW prototype with provisions to have both configurations to settle the issue.  K1 and effective lengths were calculated for all dipole cases and fed back to David Douglas for inclusion in his engineering version of the lattice as well as the final physical iron lengths on the magnet drawings. The gap in the 180° bend was changed to 3 inches from 2 inches to reflect the thinking that beam tube dimensional changes cause wake fields that are unacceptable.

Quadrupoles
Prototype Fabrication: Inspection of the aluminum sample quadrant showed conflicting results between CMM and optical tooling measurements of 0.005" and 0.010".  We decided to concentrate on building an assembled core.  This will get us a prototype sooner. An inspection will be performed of the assembled quadrants and any required "corrective machining" will be performed on the assembly.  By months end the four cores were received and we started inspection.  Three "affordable" quotes were received for machining cores.  We are awaiting three additional quotes from coil vendors.  An encouraging quote was received from Master Machine who proposed a mix of wire EDM and CNC machining.

Measurement Probe: Glass and plastic material for the probe arrived.  The 50-turn litz wire coil was completed and fabrication of the 100-turn continues to be built in-house by a technician in the HPEE Group.  Drawings for the probe body are in the final steps of detailing.

WBS 10 (Wiggler):

Drawings for the manifold mounts were revised to accommodate mounting the wiggler cover and are in procurement.  Drawings for the buss bar mounting were completed and are also in procurement.

The dispersion section vendor has all the materials for the magnet and has started fabrication. A visit to the vendor is planned for the beginning of January The drawings for the wiggler viewers have been completed and signed off.  A mistake was found in the optics calculations for the dispersion section viewer.  A new solution was found which allows us to use all the same parts for the dispersion section viewer and the wiggler viewers.   This reduces the number of drawings necessary.
The 3D design of the wiggler vacuum chamber support was finalized. The modified pole clamps have been attached  to one of the wigglers.   Inspection of the gap showed a maximum spread of 0.0025" (spec 0.002".)   This shows that the design is acceptable.

The supplier of Group-3 Hall Probes (GMW) was contacted regarding information on planar Hall effect specifications.  Specifications for magnetic measurements were also sent to STI Optronics for their review and quotation.

WBS 11 (Optics):

We received bids for the mirror test stand vacuum vessel and the work has been awarded to a local vendor.  All the optical windows for the test stand have been ordered. We have ordered a gimbal mount that will be used in the test stand, and is being considered for the optical cavity as well.  We will use it in our design of the optical cavity, so we can design a vacuum vessel for it, and feed that information back to WBS 3 in order for them to more quickly finalize the lattice.  We are working with the design engineer at AES responsible for finalizing the design for the HR mirror.

We received the quad pyros and associated electronics from the I & C group, and this was installed onto the O-BPM mechanical assembly in the optics control room (OCR). It has resulted in a much improved ability to determine and fix the beam position between the OCR and ModeMaster at the end-of-line (EOL ).
The custom-built enclosure for the OCR optical table that allows more efficient purging was installed last week.

Operations/Commissioning:

Using a new outcoupler optimized for 1 micron operation, we measured 220 W of power at 1028 nm upstairs, and 340 W downstairs, almost immediately after the outcoupler.  These results solidly place us in record-making levels of power for an ultrafast laser.  The ratio 65%, agrees with a calculation based on the literature values of silver's reflectivity.  We took this light (while in pulsed mode), and doubled it in lithium iodate, measuring a conversion efficiency of ~ 28% to ~ 530 nm.  This efficiency dropped considerably when we attempted to double while running cw, due to crystal heating.  Then, working with W. Cooke (College of William and Mary), we tripled and quadrupled the ~ 1 micron light.  The quadrupling conversion efficiency was actually fairly reasonable, ~ 20% from the SHG light, or ~ 5% from the ~1 micron light.  The third harmonic generation conversion efficiency was poor, due to temporal dispersion effects in the material.  We're obtaining more optimized materials, but do not expect them until later this month.  All this work is important, both for our current users wishing to do micromachining, or AMO physics.

We also delivered beam to a group from Kansai Research Institute and RPI, here to generate high intensities of THz radiation.  To date, they have been very successful, measuring both relatively high intensities, and the duration of the radiation.

During the  last week of FEL operations for the Fall User run  (Nov. 13-17), the FEL ran well and very stably all week. Operations were devoted to providing beam to G. Luepke (College of William and Mary) to study the dynamics of H-implanted defects in Si.  It pushes the performance of the IR Demo to the limits in terms of amplitude, wavelength, and position stability.  So far, we've been able to meet these requirements.
 




Privacy and Security Notice